OpenStudy (lgbasallote):

obtain a family of solutions for \[(x-2y)dx + (2x+y)dy = 0\]

OpenStudy (anonymous):

put it in this form first: dy/dx=-x+2y/2x+y then devide by x the right side dy/dx=-1+2(y/x)/2+(y/x) make variable change y/x=u y=ux dy/dx=u+x(du/dx) substitute in the DE and solve by separating variables

OpenStudy (lgbasallote):

why cant you sub from the start?

OpenStudy (anonymous):

you could, just is more clear other way

OpenStudy (lgbasallote):

i see

OpenStudy (anonymous):

since you don't mess with differentials

OpenStudy (lgbasallote):

is u + x(du/dx) = udx + xdu?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

great..thanks